OpenStudy (anonymous):

Please help me? :) Rationalize the denominator of √-49/(7-2i) - (4+9i) My answer: 49+21i/58

OpenStudy (anonymous):

plz put brackets after each expression so that ur ques get clear to me

OpenStudy (anonymous):

ok koala ;0

OpenStudy (anonymous):

√-49/(7-2i) - (4+9i) is this like this? \[\sqrt{(\frac{ -49 }{ 7-2i })}-(4+9i)?\]

OpenStudy (anonymous):

or is it like this? \[\sqrt{\frac{ -49 }{ 7-2i }-(4+9i)}?\]

OpenStudy (anonymous):

yes!!! :))

OpenStudy (anonymous):

or is it like this? \[\sqrt{\frac{ -49 }{(7-2i)-(4+9i) }}\]?

OpenStudy (anonymous):

which one now? the first, second or the third?

OpenStudy (anonymous):

the former was correct ;)

OpenStudy (anonymous):

the first one :)

OpenStudy (anonymous):

ok the first one is \[\sqrt{\frac{ -49 }{ 7-2i }( \frac{ 7+2i }{ 7+2i })}-(4+9i)\] \[\sqrt{\frac{ -49(7+2i) }{ 49-4i ^{2} }}-(4+9i)\] \[\sqrt{\frac{ -49(7+2i) }{ 49-4(-1) }}-(4+9i)\] can you further simplify them ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

= sqrt[(-343-98i)/53]-(4+9i) =sqrt[-6.472-1.85i]-(4+9i)

OpenStudy (anonymous):

ok i think we cant simplify them any more :D

OpenStudy (anonymous):

thank you kind sir!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!11111

OpenStudy (anonymous):

ok YW have fun now and good luck ... :D